OpenStudy (daniel.ohearn1):

Can you explain to me how to go about integrating a product of three functions ? For example: Integrate (e^x)(ln(x))(sinx)dx

OpenStudy (irishboy123):

i am not sure that is the best example because i think you can just write it as \(\mathcal{Im} \int ~ e^x \ln(x) e^{ix} ~dx \) but if you take the product rule: \((u v w)' = u' v w + u v' w + uv w'\) perhaps you can then use \(\int u' v w ~ dx= uvw - \int u v' w ~ dx - \int uv w' ~ dx\) but without actually trying it out, i think you'd need to luck out and have two of the functions that are easy to differentiate/integrate, .....

OpenStudy (irishboy123):

@freckles

OpenStudy (daniel.ohearn1):

How did you get that formula with the integral?

OpenStudy (irishboy123):

which one?

OpenStudy (daniel.ohearn1):

uvw one